### v1

# Clear Minds Think Alike: What Makes LLM Fine-tuning Robust? A Study of Token Perplexity

Abstract
Maintaining consistent model performance across domains is a fundamental challenge in machine learning. While recent work has explored using LLM-generated data for fine-tuning, its impact on cross-domain generalization remains poorly understood. In this paper, we present a systematic analysis revealing that fine-tuning with LLM-generated data not only improves target task performance but also reduces out-of-domain (OOD) degradation compared to fine-tuning with ground truth data. Through analyzing the data sequence in tasks of various domains, we demonstrate that this enhanced OOD robustness stems from a reduced prevalence of high perplexity tokens in LLM-generated sequences. Following this hypothesis we showed that masking high perplexity tokens in ground truth training data also achieves similar OOD preservation comparable to using LLM-generated data. Extensive experiments across diverse model architectures and scales, including Gemma2-2B, Mistral-7B and Llama3-8B, corroborate the consistency of our findings. To the best of our knowledge, this work provides the first mechanistic explanation for the superior OOD robustness conferred by LLM-generated training data, offering valuable insights for developing more robust fine-tuning strategies.
1 Introduction
Supervised fine-tuning (SFT) on Large Language Models (LLMs), particularly for instruction following [20], have demonstrated remarkable capabilities in comprehending novel instructions [31] and generating useful outputs across various tasks, including summarization and web queries [26, 18, 1, 12]. As LLM applications continue to expand into specialized domains such as data querying and programming assistance [28, 19], practitioners face significant computational and data constraints when adapting LLMs. This challenge is further compounded by the performance saturation observed in smaller LLMs, where benchmark scores on tasks like MATH [10] and MBPP[2] plateau under limited parameter sizes, probably due to inherent limitations in model capacity. Fine-tuning such models often risks degrading their general capabilities [23, 16, 22]. A critical challenge emerges: how to efficiently fine-tune pre-trained instruction-following LLMs for domain-specific tasks while preserving their general capabilities, especially when the original fine-tuning dataset is inaccessible or when computational resources are limited.
[…]
Figure 1: An example MATH problem showing high perplexity tokens (highlighted in red, perplexity $\geq 2.5$) in both ground truth and Self-Output responses.
Recent work has demonstrated the effectiveness of using model-generated data for fine-tuning, showing that training on high-quality samples from model generations [36, 8, 24] can outperform training on ground truth data. While Ren et al. [23] showed that model-generated responses yield superior performance compared to human annotations, their analysis focused on using more capable models (e.g. GPT-4) to generate training data and relieve the degradation of general tasks. However, although they mange to use "familiarity", or low perplexity, of generated response to explain how unfamiliar data causes lower performance, it is still not fully understood that how generated data works well for instruction tuning: (1) Context of regenerated response differs from ground truth response so the bias in context could disturb the effectiveness of low perplexity training in their experiments. (2) the impact of low perplexity training on out-of-domain performance is still poorly understood, and (3) The distillation of larger-sized model generally benefits the training, it is better to eliminate the effect of larger-sized model and discuss the low perplexity training issue with the same models. Our work provides the first mechanistic explanation for why LLM-generated data leads to more robust fine-tuning outcomes. Through systematic analysis, first, we demonstrate that fine-tuning with LLM-generated data significantly reduces out-of-domain degradation compared to ground truth data training. Second, we trace this improvement to a key characteristic as Figure 1 shows. The LLM generated response has lower sentence perplexity and lower proportion of high perplexity tokens. Hence, this finding leads to a practical insight: simply masking high perplexity tokens in ground truth training data achieves comparable out-of-domain preservation to using LLM-generated data for training. In such way, we also manage to use the same model before training to eliminate high perplexity tokens from the original ground truth context, and discover the improvement that benefits from low perplexity training.
Our contributions are threefold:
  • We identify the mechanistic explanation for out-of-domain performance degradation through our proposed method: Selective Token Masking (STM), a simple yet effective training strategy to filter out high perplexity tokens to achieve low perplexity training.
  • We provide the first empirical evidence that LLM-generated training data significantly reduces out-of-domain degradation during fine-tuning.
  • We demonstrate the broad applicability of these findings through comprehensive experiments across different model architectures and scales, including Gemma2-2B, Mistral-7B, and Llama3-8B.
These insights offer new directions for developing robust fine-tuning strategies that better preserve general capabilities while enhancing target task performance.
2 Training LLMs on Self-Generated Data
To examine the effectiveness of LLM-generated data for fine-tuning, we investigate two distinct methods for generating training data of two different tasks and evaluate the models trained with the generated data on five different tasks. Together, these two methods, Self-Output and Rephrase, provide a framework for generating LLM-based training data, each addressing different challenges and trade-offs. In this section, we outline our original datasets for LLMs to generate from, and the approaches to creating self-generated training datasets with language models.
2.1 Training and Evaluation Framework
For training in-domain task, we adopt MBPP and MATH datasets for training. These two datasets provide complete guidance and comprehensive ground truth labels for models to learn from, including complete solutions, reasoning paths, and test cases, not just final answers. For evaluation, besides the testing set of MBPP and MATH, we also evaluate on GSM8k [5], ARC-Challenge [4] and BIRD [14] to evaluate fine-tuned models performance in different aspects of generation. Detailed of dataset will be introduced in Section 4.
2.2 Self-Output
Self-Output method follows the same high level concept as other synthetic generation works such as [36, 23] where we start by sampling a diverse LLM response via strict filtering process thereby creating a high-quality synthetic dataset. For each training instance, we use a language model $M$ such as Llama3-8B Instruct to generate $N$ (e.g. 32) distinct responses with temperature $T=0.7$. We then filter the generated responses which are semantically aligned with the ground truth to ensure the quality of our synthetic dataset.
A key advantage of Self-Output is its ability to produce reliable training data when the ground truth is verifiable, such as in programming tasks or other domains with unambiguous correct answers. However, this method is limited to scenarios where such verification is feasible, as it relies on the availability of a clear ground truth for filtering.
Table 2 presents the performance comparison between Self-Output, Rephrase, Ground Truth training and base models across both in-domain and out-of-domain tasks using different models in terms of model size and model series. The base models in this paper refer to generic instruction following models before training, e.g. Llama3-8B-Instruct [6], Gemma2-2B Instruct [27] and Mistral-7B Instruct v0.3 [11]. For in-domain performance, Self-Output demonstrates consistent improvements in both MBPP and MATH compared to their respective Ground Truth baselines. For out-of-domain scenarios, the Self-Output method achieves robust performance compared with the base models in most cases, and some could improve. We also observe a particularly strong transfer from MATH training to GSM8K evaluation, though performance on the BIRD dataset remains challenging across all methods. These results suggest that our Self-Output method enhances model generalization, particularly in mathematical reasoning tasks, while maintaining competitive performance in programming and general knowledge domains. For more details, such as performance comparison with other different models, please refer to Appendix A.
2.3 Rephrase
Yang et al. [35] proposed self-distillation. The process uses an instruction-finetuned LLM to rephrase ground-truth responses in its own style. By providing both the instruction and ground truth to the target LLM, it generates semantically equivalent reformulations with ground truth response. Compared to Self-Output, Rephrase is computationally efficient, as it requires only a single pass of generation without the need for additional filtering. However, Rephrase is susceptible to hallucinations, as evidenced by the higher variance in token level perplexity (see Table 1). Despite this limitation, Rephrase offers versatility, as it does not rely on verifiable ground truth and can be applied to a broader range of tasks. We apply Rephrase as one of the baselines in Table 2.
Together, these methods provide a framework for generating LLM-based training data, each addressing different challenges and trade-offs. In the following section, we analyze the perplexity of datasets generated by these methods to quantify their impact on model uncertainty and fine-tuning stability. These insights offer new directions for developing robust fine-tuning strategies that better preserve general capabilities while enhancing target task performance.
2.4 Analysis of Perplexity Across Each Samples
In this section, we analyze the perplexity of datasets generated by Self-Output and Rephrase to quantify their impact on model uncertainty and fine-tuning stability. These insights offer new directions for developing robust fine-tuning strategies that better preserve general capabilities while enhancing target task performance. We first collect the Llama3-8B Instruct inference statistics from all three methods: Ground Truth, Rephrase and Self-Output strategy. We evaluate the perplexity of responses by summing up the negative log probability of each token conditioned on its previous context in the sentence to measure the uncertainty of model outputs over the entire sequence. Specifically, the perplexity $PPL$ [^1] In this section we use natural log instead of base 2 for computation convenient across a sequence is computed as:
$PPL=\exp(\frac{1}{N}\sum_{i=1}^{N}P(w_{i}|w_{1},...,w_{i-1}))$ (1)
where $P(w_{i}|w_{1},...,w_{i-1})$ represents the conditional probability of token $w_{i}$ given its preceding context, and $N$ is the total number of tokens in the sequence.
Table 1: Perplexity average and variance score over answer sequences of MBPP and MATH training datasets on Llama3-8B Instruct. The average sentence perplexity are calculated by averaging the sum of each sentence’s perplexity over token perplexity score.
Data Method Average Perplexity
MBPP Ground Truth 4.83 (7.04)
Rephrase 1.69 (0.16)
Self-Output 1.16 (0.01)
MATH Ground Truth 2.45 (0.81)
Rephrase 2.11 (9.28)
Self-Output 1.34 (0.03)
Table 1 demonstrates that the Self-Output strategy achieves both the lowest average perplexity of 1.16. Among the three approaches, Ground Truth exhibits the highest uncertainty with mean perplexity of 4.83, and the perplexity of Rephrase are slightly higher than Self-Output. The consistently lower uncertainty metrics of Self-Output suggest that LLMs are less likely to deviate far from their initial weights when fine-tuning on downstream tasks created by Self-Output. This characteristic maintains model original performance on out-of-domain tasks while effectively learning from the in-domain task training data.
Figure 2: Comparison of token-level perplexity (PPL) distributions between human-annotated ground truth (left) and Llama3-8B generations for Rephrase and Self-Output sequence (middle and right).
To investigate token-level predictability we sample a single MATH response and visualize the token level perplexity value. Figure 2 reveals distinct patterns across Ground Truth, Rephrase, and Self-Output responses. The first 20 tokens variance are 2.04 for Ground Truth, 2.31 for Rephrase, and 0.04 for Self-Output which confirm our observation: Ground Truth exhibits frequent perplexity spikes of high-amplitude (perplexity around 6-8), which reflects the natural variability in human problem-solving expressions. While Rephrase maintains similar peak amplitudes to Ground Truth, i.e. more regular spike patterns, it suggests a hybrid variability between human and LLMs. In contrast, Self-Output demonstrates remarkably consistent low token perplexity values rarely exceeding 2, indicating highly predictable token sequences. These patterns suggest that Self-Output generates more deterministic solutions, potentially beneficial for maintaining consistent problem-solving approaches, while human responses introduce natural variations.
To demonstrate the different content of tokens distribution in between Ground Truth and Self-output responses, please refer to Appendix B.
3 Selective Token Masking (STM)
Previous analysis of Figure 2 reveals a significant discrepancy in token perplexity distribution between Self-Output generated data and other approaches. This observation raises a fundamental question: Is the superior performance of Self-Output methods directly attributable to the lower perplexity of their generated tokens? To investigate this hypothesis and develop more efficient fine-tuning strategies, we propose Selective Token Masking (STM), a novel approach for supervised fine-tuning (SFT) that focuses on the role of token perplexity in model training.
3.1 Selective Token Masking
The core mechanism of STM is straightforward yet effective: using an existing instruction-tuned model to compute token perplexities, we mask and exclude ground truth tokens that exceed a predefined threshold $\tau$ during the fine-tuning process. We investigate the foundations of successful fine-tuning through two key findings. First, STM reveals that token perplexity plays a crucial determinant of fine-tuning outcomes. Second, STM demonstrates that the performance gains of Self-Output methods can be achieved through direct perplexity-based token management, suggesting these gains stem from low perplexity token distributions rather than the self-generation process itself. This approach not only provides computational efficiency but also establishes a systematic framework for understanding model adaptation mechanisms.
3.2 Comparison
Our approach shares some conceptual similarities with [15], yet differs significantly in implementation and efficiency. Their method requires a two-stage process: first training a reference model on high quality data to identify high perplexity tokens from ground truth data as learnable tokens, and then pre-trains an LLM on the learnable tokens. In contrast, STM achieves similar objectives through a streamlined single-stage process: it directly utilizes an existing instruction-tuned model to compute token perplexities via a single forward pass. Thus, STM eliminate the computational overhead of [15], while sustaining to the performance of low perplexity training.
4 Experiment Setting
To evaluate the effectiveness of our proposed method, we select benchmark datasets with train sets designed to evaluate the robustness on the in-domain and out-of-domain data after fine-tuning. We only chose datasets with a verifiable ground truth such that we can filter incorrect self-generated responses out for better data quality.
Table 2: Performance for Ground Truth, Rephrase, Self-Output, and STM methods trained on datasets MBPP and MATH respectively. Values of a cell block in gray represent the in-domain testing performance of the model trained with in-domain task data.
Model in-domain method MBPP MATH ARC gsm8k BIRD
task
Gemma2-2B - base model 51.06 28.10 74.74 58.45 13.89
MBPP ground truth 39.95 $\downarrow$ 22.90 $\downarrow$ 71.59 $\downarrow$ 19.26 $\downarrow$ 5.14 $\downarrow$
Rephrase 48.41 $\downarrow$ 28.90 $\uparrow$ 70.56 $\downarrow$ 48.22 $\downarrow$ 13.75 $\downarrow$
Self-Output 53.44 $\uparrow$ 29.20 $\uparrow$ 58.45 $\downarrow$ 52.39 $\downarrow$ 13.95 $\uparrow$
$STM_{\tau=2.5}$ 49.21 $\downarrow$ 29.80 $\uparrow$ 74.83 $\uparrow$ 56.41 $\downarrow$ 13.75 $\downarrow$
MATH ground truth 45.77 $\downarrow$ 21.70 $\downarrow$ 29.52 $\downarrow$ 19.03 $\downarrow$ 12.45 $\downarrow$
Rephrase 46.83 $\downarrow$ 20.00 $\downarrow$ 73.38 $\downarrow$ 37.00 $\downarrow$ 12.58 $\downarrow$
Self-Output 50.00 $\downarrow$ 30.90 $\uparrow$ 72.95 $\downarrow$ 57.01 $\downarrow$ 13.89 -
$STM_{\tau=2.5}$ 49.47 $\downarrow$ 30.30 $\uparrow$ 75.34 $\uparrow$ 55.42 $\downarrow$ 13.30 $\downarrow$
Llama3-8B - base model 59.26 31.40 78.16 72.55 20.53
MBPP ground truth 57.94 $\downarrow$ 23.30 $\downarrow$ 59.13 $\downarrow$ 18.50 $\downarrow$ 17.60 $\downarrow$
Rephrase 54.06 $\downarrow$ 30.00 $\downarrow$ 75.43 $\downarrow$ 72.40 $\downarrow$ 19.16 $\downarrow$
Self-Output 61.64 $\uparrow$ 33.40 $\uparrow$ 79.18 $\uparrow$ 78.39 $\uparrow$ 20.01 $\downarrow$
$STM_{\tau=2.5}$ 61.11 $\uparrow$ 32.00 $\uparrow$ 78.92 $\uparrow$ 75.43 $\uparrow$ 20.66 $\uparrow$
MATH ground truth 58.73 $\downarrow$ 25.80 $\downarrow$ 48.63 $\downarrow$ 67.93 $\downarrow$ 19.42 $\downarrow$
Rephrase 60.32 $\uparrow$ 32.90 $\uparrow$ 78.50 $\uparrow$ 78.62 $\uparrow$ 18.90 $\downarrow$
Self-Output 60.05 $\uparrow$ 34.40 $\uparrow$ 80.12 $\uparrow$ 78.01 $\uparrow$ 19.88 $\downarrow$
$STM_{\tau=2.5}$ 61.64 $\uparrow$ 33.40 $\uparrow$ 75.09 $\downarrow$ 75.06 $\uparrow$ 20.27 $\downarrow$
4.1 In Domain and Out Domain Datasets
We focus on three domains—Programming, Mathematics, and Knowledge-based tasks—to study performance shifts after supervised fine-tuning (SFT) with LoRA on each domain. For the Programming and Mathematics domains, we select MBPP and MATH for training, while others and Knowledge-based tasks for out-of-domain evaluation to assess generalization capabilities.
4.1.1 Programming Tasks
For code generation evaluation, we use MBPP for both training and evaluation, while BIRD serves as an out-of-domain assessment. This pairing tests both direct programming ability and cross-language generalization from Python to SQL.
MBPP: A Python programming benchmark containing 974 problem-solution pairs. We partition training set into 374 train, 90 validation and using the original 378 test examples. Performance is evaluated using the pass@1 metric.
BIRD: A Text-to-SQL generation benchmark that requires models to translate natural language queries into executable SQL statements. Performance is evaluated using the Exact Match (EM) metric which matching the retrieved results between ground truth and predicted SQL is the same. We use this dataset exclusively for evaluation.
4.1.2 Mathematical Reasoning Tasks
We use MATH for training and evaluation, and GSM8K for out-of-domain testing. These datasets differ in format: MATH features competition-style problems, while GSM8K uses natural language, allowing us to assess generalization across mathematical expression styles.
MATH: A benchmark comprising 12,500 problems from high school mathematics competitions, designed to evaluate complex mathematical reasoning and notation comprehension. We use this dataset for both training and evaluation.
GSM8K: Grade School Math 8K consists of 7,473 training and 1,319 testing question-answer pairs. We utilize only the test set for out-of-domain evaluation, leveraging its natural language format to assess generalization from formal to informal mathematical reasoning.
4.1.3 Knowledge-based Task
To assess broader generalization capabilities beyond mathematical and programming domains, we incorporate:
ARC-Challenge: A specialized subset of the ARC science question dataset, comprising 2,590 questions (1,119 training, 299 validation, and 1,172 test) selected for their increased difficulty and reduced susceptibility to statistical shortcuts. This dataset is used for evaluation.
4.2 Models
We evaluate our method across different model scales and architectures using three publicly available instruction-tuned language models; here they are referred to as Base Models:
  • Gemma2-2B Instruct: A 2.6 billion parameter model designed for resource-efficient deployment
  • Mistral-7B v0.3: A 7 billion parameter model featuring grouped-query attention and sliding window attention mechanisms.
  • Llama3-8B Instruct: An 8 billion parameter model which has the best overall performance in all 3 models.
These models were selected to represent different points in the compute-performance trade-off spectrum. Llama3-8B and Mistral-7B represent mid-scale models with different architectural innovations, while Gemma2-2B allows us to assess our method’s effectiveness on more resource-constrained settings. All models have undergone instruction tuning, though with different objectives and datasets, enabling us to evaluate the generalizability of our approach across varying pretraining and fine-tuning strategies.
5 STM Experiement Results
Table 2 shows that by setting a threshold to filter out high perplexity tokens as 2.5, which approximately filtering out around 20% to 24% of total tokens, with original instruct model on the ground truth dataset as training data, the testing performance is comparable to the Self-Output method. This strategy not only maintains in-domain performance but also preserves out-of-domain generalization. This result agrees with our hypothesis that the presence of high perplexity tokens is one of the causes of performance degradation of ground truth model, and low perplexity tokens sufficiently mitigate such negative impact for strong generalization without the need for self-generated training data.
Table 3: STM performance comparison between Gemma2-2B Instruct and Llama3-8B Instruct. The percentage indicates the number of filtered tokens from the training data.
Model MATH MBPP GSM8K ARC BIRD
Gemma2-2B Instruct
Ground Truth 21.7 45.8 19.0 29.5 5.1
$STM_{\tau=1000}$ (2.5%) 23.3 48.9 24.9 25.4 12.8
$STM_{\tau=25}$ (9.3%) 27.0 48.2 49.7 72.9 13.1
$STM_{\tau=2.5}$ (23.8%) 30.3 49.5 55.4 75.3 13.3
Llama3-8B Instruct
Ground Truth 25.8 58.7 48.6 67.9 19.4
$STM_{\tau=1000}$ (1.7%) 27.7 59.2 69.3 71.2 19.4
$STM_{\tau=10}$ (11.0%) 33.0 61.1 77.6 75.6 19.5
$STM_{\tau=2.5}$ (22.2%) 33.4 61.6 75.1 75.1 20.3
5.1 Optimal threshold selection for STM
A key hyperparameter of STM is the perplexity threshold to filter tokens from the supervised fine-tuned (SFT) model. We explore the existence of an optimal threshold that maximizes performance. Figure 3 reveals that filtering approximately 24% of tokens yields the best results. This optimal threshold demonstrates consistent benefits across both in-domain and out-of-domain datasets, including GSM8K, ARC, and MBPP, showing its robustness and generalizability similar to Self-Output data.
Figure 3: MATH SFT using STM method on Llama3-8B Instruct on different levels of token filtering levels. The best in domain performance nearly matches peak performance on out of domain tasks : GSM8k, ARC, MBPP as well.
5.2 Alternative Token Selection Strategies
In addition to that STM uses initial model perplexity for token selection, we investigate an alternative strategy.
Cross-scale Filtering: We explore perplexity assessment using larger models within the same model family (e.g., Gemma2-9B Instruct) to guide token selection for smaller models (e.g., Gemma2-2B Instruct), as denoted as ${STM}_{9B\tau=2.5}$. This approach investigates whether token selection benefits from a more robust understanding of larger models, potentially offering a form of knowledge distillation through perplexity-based filtering.
Table 12 shows the results for the best-performance models of the original STM and the alternative of STM. Compared to base model performance, both STM and $STM_{9B}$ outperform the base model results in term of in-domain and out-of-domain tasks, showing STM’s effectiveness in larger scaled models still benefits model training. However, even though selecting similar filter rate, scaling up perplexity model does not outperform the original STM. It is likely that a more capable model could score with lower perplexity on tokens that smaller-sized models perceive as high perplexity tokens. Therefore, to optimize the performance of STM, our token selection strategy is preferable to apply the same model as the training model, and set a threshold to match an ideal filtering rate(around 20% to 25%). We also investigated other alternatives, please see Appendix D.
Table 4: List of alternative token selection methodology supervised fine-tuned on MATH in Gemma2-2B Instruct. Here we choose Gemma2 series because Gemma2 has feasible model sizes to implement the cross-scale setting for comparison.
Gemma2-2B Filter MATH MBPP GSM8K ARC BIRD
Base model 0% 28.1 51.1 58.5 74.7 13.9
Ground Truth 0% 21.7 45.8 19.0 29.5 5.1
$STM$ 23.8% 30.3 49.5 55.4 75.3 13.3
${\tau=2.5}$
${STM}_{9B}$ 21.4% 27.2 47.6 42.0 72.4 14.1
${\tau=2.0}$
6 Additional Analysis Results
Other than STM experiment results, we conducted additional analysis on the learning and robustness of our mentioned methods.
6.1 Does Self-Output Response Correctness Matters?
In our experiment, the lowest perplexity scored data are collected exclusively from correct Self-Output subsets. In MBPP training task, a correct Self-Output data refers to a generated response that passes all the test cases. In MATH, correctness refers to a correct final answer parsed from a response. Hence, it raises an important research question regarding the necessity of correctness criteria when simultaneously pursuing in-domain performance improvements and minimizing out-of-domain degradation. In Table 5, we collect different correctness rates of Self-Output response as MBPP training datasets respectively with Llama3-8B Instruct model. Correctness rate refers to the ratio of the samples whose labels are verified as "correct". Our results show that the correctness of MBPP training data strongly affects the in-domain performance. It is intuitive that training with incorrect answers leads to lower in-domain performance. However, compared to model trained on ground truth data (the correctness rate is 100%), which fails to sustain out-of-domain task performance, models trained with higher correctness rate of Self-Output data achieves higher in-domain improvements and still sustains the out-of-domain task performance robustness at the same time. Therefore, high correctness improves in-domain performance, while Self-Output data effects out-of-domain performance more than correctness. We also have the correctness test with smaller models like Gemma2-2B Instruct, please refer to Appendix E.
Table 5: The performance of models trained with different correctness rate of MBPP Self-Output dataset and trainable tokens number. Values of a cell block in gray represents the in-domain testing performance of the model trained with in-domain task data.
rejection rate Trained tokens MBPP MATH ARC GSM8k BIRD
GT 75,464 .579 .233 .591 .185 .176
Self-output
100% 65,602 .606 .324 .793 .773 .199
75% 69,605 .593 .346 .794 .778 .199
0% 59,738 .561 .315 .802 .770 .188
base model - .593 .314 .782 .726 .206
6.2 Convergence Performance Across Different Methods
Training low perplexity data also leads to a low training loss intuitively, and model could converge within fewer training epochs. Figure 4 showcases the difference of training procedure on the ground truth data and the Self-Output data on MBPP task. The results show that training with Self-Output data converges as a lower validation loss curves within half of the training steps (3 epochs v.s. 7 epochs) compared to training with ground truth data. Moreover, within small steps of training (3 epochs), there is a higher testing performance of Self-Output data training (improved $+1.41\%$ compared to original Llama) than the performance of ground truth data training (degraded $-6.62\%$). STM method also shows the similar trends. This analysis suggests that it is more efficient to train with low perplexity data to achieve higher in-domain performance than to train with high perplexity data.
Figure 4: MBPP in-domain testing accuracy, validation and training loss of ground truth, Self-Output, and STM data with 2.5 as perplexity filtering threshold.
6.3 Can Regularization Methods Train Robust Models?
We examine alternative regularization techniques for mitigating out-of-domain performance degradation during in-domain fine-tuning. Our comparative analysis encompasses two widely-adopted regularization approaches: Dropout [25] and Weight Decay [13]. These methods have demonstrated effectiveness in preventing model from overfiting. Table 6 shows that tuning weight decay and dropout rate is not a robust approach to achieving in-domain performance improvement and avoiding out-of-domain performance degradation at the same time, while Self-Output data training simply improves both. For more detail about the experiment settings and results, we leave them in appendix F.
Table 6: Comparison between Self-output model, Base model and Ground Truth models applying different combinations of regularization parameters. *Note that the regularization parameters of the highest MBPP and MATH performances are not the same.
Model Self Base Regular-
finetuned on MBPP Out Model ization
MBPP .616 .598 .558*
MATH .334 .314 .329*
6.4 LoRA Weight Analysis
We also investigate the relationship between token level perplexity and the characteristics of weight updates in Low-Rank Adaptation (LoRA). Our analysis focuses on understanding how different types of training data affect the magnitude and dimensionality of parameter adjustments in transformer-based language models. Given that LoRA effectively learns offset weights $\Delta W$ to the original model parameters $W$, we calculate the L2 norm of $\Delta W$ to quantify the degree of deviation introduced by each fine-tuning dataset. As shown in Table 7, fine-tuning with Self-Output data consistently results in smaller L2 norm of weight updates compared to Rephrase and Ground Truth data, while STM produces the smallest updates (0.45-0.56) across all models. These minimal perturbations suggest that both approaches, particularly STM, help preserve the model’s original capabilities.
Table 7: L2 norm of $\Delta W$ larger value signifies a larger update from the original weight $W$
Model Self Re- Ground STM
finetuned on MBPP Out phrase Truth
Llama3-8B Instruct 6.53 7.31 17.75 0.55
Gemma2-2B Instruct 4.03 5.78 5.69 0.45
7 Related Work
Instruction-tuning
Instruction following [21, 3] is widely used to align the responses of LLMs with target task values. Practically, instruction following requires instruction tasks dataset to fine-tune LLMs, where each dataset consists of instructions and desired responses. Hence, the performance of instruction-tuning heavily relies on the quality of instruction data such as context richness [33, 29, 37]. In addition, Ghosh et al. [7] investigated the limitation of instruction-tuning of LLMs about the performance degradation (as known as catastrophic forgetting) from pattern-copying behaviors and hallucinations with Lora/full fine-tuning. These works give us a good start on how performance degrades in terms of models’ response behaviors and benchmarks when fine-tuning with instruction following dataset.
Using LLM-generated data for instruction fine-tuning
As [29, 3] shows, since LLM breakdown easily after training with different tasks. There are several remedies to the performance improvement focusing on training data augmentation [30, 34]. For instance, [23] use much larger LLM such as GPT-4 and Claude to generate responses of questions as training labels, which improves both the target task and others non-target tasks performances. However, such distillation method neglects the correctness of generated labels for training data so that incorrect responses could be generated as the amount of generated data increases, and using equal or smaller size models for distillation could be challenging. [35] prompts LLMs to simply rephrase response of existing ground truth to generate labels to match similar styles of the LLMs for fine-tuning. However, rephrasing ground truth answer limit the output distribution and result in lower performance in our study. Furthermore, [9] exploits a base LLM as a judge to pick out answerable questions and unanswerable questions to compose a new training dataset to improve in-domain performance only on QA/conversation dataset. Although using Mistral-7b v0.2 to generate acceptable responses for answerable question improves in-domain and out-of-domain tasks, it is rarely discussed that if the proposed method is applicable for different model sizes and series. Besides, using LLM as a judge still brings noise to the correctness of training, which shows to be important for in-domain training in our study. Lastly, [15, 17] propose a Selective Language Modeling (SLM) to score favorable tokens from training data, and training on selected tokens could bring higher performance. However, the requirements for pre-training a reference model or prepare a high-quality dataset for scoring model could be exhausting depending on task difficulty. Our proposed STM does not require such training before selection of tokens, which brings both efficiency and performance improvement at the same time.
8 Conclusion
In this work, we present a possible explanation for the effectiveness of fine-tuning on LLM outputs. Our analysis and empirical result reveal that the improved robustness stems from reduced high perplexity tokens due to Self-Output sequences. We proposed an alternative method, STM, which filters out high perplexity tokens from ground truth data. We conducted extensive experiments across Gemma2-2B, Mistral-7B and Llama3-8B, and the results showed that the STM method is competitive with Self-Output data in both in-domain and out-of-domain tasks. Our findings suggest that token-level perplexity plays an important role in the fine-tuning process, offering future robust fine-tuning a simple but effective baseline.
[…]
Appendices
Appendix A Other LLM series performance
Similar to Table 2, we also have performances of other models. Due to space limitation, we add the results of Mistral-7B Instruct model of Self-Output, Rephrase, Ground Truth, and STM performances here as Table 8 shows: our conclusion that Self-Output and STM strategies for training data generation still holds for different model sizes and series.
Table 8: Performance for Ground Truth, Rephrasal, Self-Output, and STM methods across datasets MBPP and MATH. Values of a cell block in gray represent the in-domain testing performance of the model trained with in-domain task data.
Model in-domain method MBPP MATH ARC gsm8k BIRD
task
Gemma2-2B - base model 0.5106 0.2810 0.7474 0.5845 0.1389
MBPP ground truth 0.3995 $\downarrow$ 0.2290 $\downarrow$ 0.7159 $\downarrow$ 0.1926 $\downarrow$ 0.0514 $\downarrow$
rephrase 0.4841 $\downarrow$ 0.2890 $\uparrow$ 0.7056 $\downarrow$ 0.4822 $\downarrow$ 0.1375 $\downarrow$
Self-Output 0.5344 $\uparrow$ 0.2920 $\uparrow$ 0.5845 $\downarrow$ 0.5239 $\downarrow$ 0.1395 $\uparrow$
stm 0.4921 $\downarrow$ 0.2980 $\uparrow$ 0.7483 $\uparrow$ 0.5641 $\downarrow$ 0.1375 $\downarrow$
MATH ground truth 0.4577 $\downarrow$ 0.2170 $\downarrow$ 0.2952 $\downarrow$ 0.1903 $\downarrow$ 0.1245 $\downarrow$
rephrase 0.4683 $\downarrow$ $0$downarrow 0.7338 $\downarrow$ 0.3700 $\downarrow$ 0.1258 $\downarrow$
Self-Output $0$downarrow 0.3090 $\uparrow$ 0.7295 $\downarrow$ 0.5701 $\downarrow$ 0.1389 -
stm 0.4947 $\downarrow$ 0.3030 $\uparrow$ 0.7534 $\uparrow$ 0.5542 $\downarrow$ 0.1330 $\downarrow$
Mistral-7B - base model 0.4868 0.1720 0.6408 0.3169 0.1284
MBPP ground truth 0.4550 $\downarrow$ 0.1220 $\downarrow$ 0.6826 $\uparrow$ 0.3685 $\uparrow$ 0.1617 $\uparrow$
rephrase 0.4524 $\downarrow$ 0.1910 $\uparrow$ 0.6775 $\uparrow$ 0.3889 $\uparrow$ 0.1649 $\uparrow$
Self-Output 0.4709 $\downarrow$ 0.1900 $\uparrow$ 0.6903 $\uparrow$ 0.2252 $\downarrow$ 0.1375 $\uparrow$
stm 0.4788 $\downarrow$ 0.1900 $\uparrow$ 0.6101 $\uparrow$ 0.4905 $\uparrow$ 0.1239 $\downarrow$
MATH ground truth 0.4709 $\downarrow$ 0.1670 $\downarrow$ 0.3498 $\downarrow$ $0$downarrow 0.1851 $\uparrow$
rephrase 0.4974 $\uparrow$ 0.1650 $\downarrow$ 0.7355 $\uparrow$ 0.2714 $\downarrow$ 0.1799 $\uparrow$
Self-Output 0.4683 $\downarrow$ 0.1940 $\uparrow$ 0.7312 $\uparrow$ 0.2396 $\downarrow$ 0.1447 $\uparrow$
stm 0.4656 $\downarrow$ 0.1890 $\uparrow$ 0.6510 $\uparrow$ 0.4541 $\uparrow$ 0.1167 $\downarrow$
Llama3-8B - base model 0.5926 0.3140 0.7816 0.7255 0.2053
MBPP ground truth 0.5794 $\downarrow$ 0.2330 $\downarrow$ 0.5913 $\downarrow$ 0.1850 $\downarrow$ 0.1760 $\downarrow$
rephrase 0.5406 $\downarrow$ $0$downarrow 0.7543 $\downarrow$ 0.7240 $\downarrow$ 0.1916 $\downarrow$
Self-Output 0.6164 $\uparrow$ 0.3340 $\uparrow$ 0.7918 $\uparrow$ 0.7839 $\uparrow$ 0.2001 $\downarrow$
stm 0.6111 $\uparrow$ 0.3200 $\uparrow$ 0.7892 $\uparrow$ 0.7543 $\uparrow$ 0.2066 $\uparrow$
MATH ground truth 0.5873 $\downarrow$ 0.2580 $\downarrow$ 0.4863 $\downarrow$ 0.6793 $\downarrow$ 0.1942 $\downarrow$
rephrase 0.6032 $\uparrow$ 0.3290 $\uparrow$ 0.7850 $\uparrow$ 0.7862 $\uparrow$ 0.1890 $\downarrow$
Self-Output 0.6005 $\uparrow$ 0.3440 $\uparrow$ 0.8012 $\uparrow$ 0.7801 $\uparrow$ 0.1988 $\downarrow$
stm 0.6164 $\uparrow$ 0.3340 $\uparrow$ 0.7509 $\downarrow$ 0.7506 $\uparrow$ 0.2027 $\downarrow$
A.1 STM performance on different models
Table 9 shows the overall performance of STM strategy on three different models of different sized and architecture.
Table 9: STM performance comparison between Gemma2-2B Instruct, Mistral-7B-v0.3 Instruct and Llama3-8B Instruct. The percentage indicates the number of filtered tokens from the training data.
Model MATH MBPP GSM8K ARC BIRD
Gemma2-2B Instruct
Ground Truth 21.7 45.8 19.0 29.5 5.1
$STM_{\tau=1000}$ (2.5%) 23.3 48.9 24.9 25.4 12.8
$STM_{\tau=25}$ (9.3%) 27.0 48.2 49.7 72.9 13.1
$STM_{\tau=2.5}$ (23.8%) 30.3 49.5 55.4 75.3 13.3
Mistral-7B Instruct v0.3
Ground Truth 16.7 47.1 0.0 35.0 16.2
$STM_{\tau=1500}$ (3.7%) 14.9 51.6 31.0 58.0 15.2
$STM_{\tau=30}$ (10.8%) 16.1 41.6 41.6 62.0 14.0
$STM_{\tau=3}$ (21.9%) 18.9 46.6 45.4 65.1 11.7
Llama3-8B Instruct
Ground Truth 25.8 58.7 48.6 67.9 19.4
$STM_{\tau=1000}$ (1.7%) 27.7 59.2 69.3 71.2 19.4
$STM_{\tau=10}$ (11.0%) 33.0 61.1 77.6 75.6 19.5
$STM_{\tau=2.5}$ (22.2%) 33.4 61.6 75.1 75.1 20.3
A.2 Optimal threshold selection of Gemma2-2B for STM
In addition to Self-Output results, we also examine the effectiveness of STM selection strategies with Gemma Model as Figure 5 shows. Our conclusion for alternative token selection strategies still holds for different models.
Figure 5: MATH SFT using STM method on Gemma2-2B Instruct on different levels of token filtering levels. The best in domain performance also matches peak performance on out of domain tasks : GSM8k, ARC, MBPP as well.
A.3 alternative STM strategies of Llama3-8B Instruct.
We also conduct the STM strategires on Llama3-8B Instruct model (our best model series so far). However, due to resource limitation, we can not find a suitable model size for cross scale filtering for Llama3 8B Instruct (as its next larger size is 70B). As Figure 10 shows, we still have pure STM strategy as the best masking method for the best in-domain and out-domain task performance.
Table 10: List of alternative token selection methodology supervised fine-tuned on MATH in Llama3-8B Instruct. Here we only show the STM, DPF and STM+DPF version since Llama3-8B does not have a feasible model size to implement the cross-scale setting for comparison.
Llama3-8B Filter MATH MBPP GSM8K ARC BIRD
$STM$ 22.2% 33.4 61.6 75.8 76.9 20.3
${\tau=2.5}$
$DPF$ 67.1% 23.9 57.1 48.1 70.1 18.9
${\tau=0}$
${STM+DPF}$ 70.9% 17.8 55.0 46.4 58.7 18.1
${\tau=100}$
A.4 convergence curve with Gemma2-2B Instruct
To validate the model size effect in terms of model convergence, we also observe the training procedure of Gemma2-2B Instruct. As Figure 6 shows, Self-Output and STM settings have earlier convergence time step in terms of performance and validation loss. And maintain a relative low training loss curve compared to Ground Truth data training as the same trend we see in Llama3 8B’s case.
Figure 6: MBPP in-domain testing accuracy, validation and training loss of ground truth, Self-Output, and STM data for Gemma2-2B Instruct.
Appendix B High perplexity tokens filtered by STM
We have noticed that the Self-output responses consist with low surprisal values as highly predictable token sequences. It is also worthwhile to note that the contents of high perplexity tokens between Ground Truth responses and Self-ouotput responses are different. In Math dataset, we simply categorize the tokens in model responses as Numbers, Symbols and Words. Numbers refers to the tokens which are numeric words. Symbols refers to tokens that imply a non-numeric symbols or math word in latex, such as +,,,/+,-,*,/+ , - , ∗ , /. Words refers to tokens consisting of alphabets only. As Figure 7, Figure 8, and Figure 9 shows, in terms of all the models we have tested, the proportion of number tokens with high perplexity drop drastically in Self-output responses, while the overall distribution of the three categories in both Ground Truth responses and Self-Output responses are almost the same. It implies that the number tokens in Ground Truth dataset are less predictable by models and harder to learn.
Figure 7: Ratio changes of high perplexity tokens before and after STM filtering on ground-truth dataset and Self-Output dataset of MATH task with Llama3-8B-Instruct
Figure 8: Ratio changes of high perplexity tokens before and after STM filtering on ground-truth dataset and Self-Output dataset of MATH task with Mistrak-7B
Figure 9: Ratio changes of high perplexity tokens before and after STM filtering on ground-truth dataset and Self-Output dataset of MATH task with Gemma2-2B-it
Appendix C Reproducibility
C.1 Training data cost
As we know there is efficiency difference of the data preparation between Self-Output, Rephrase and Selective Token Masking/Ground Truth. To brief about the difference. Self-Output is the most resource-exhausted since it requires to generate $N$ samples (usually set 32) and validate the correctness of samples (including parsing out the answer, and match with gold label or pass test cases), then do forward pass on those samples to pick the lowest perplexity data. On the other hand, STM only require a single forward pass on Ground Truth data to calculate perplexity of each token to do the masking in training peoxwaa. As Table LABEL:tab:efficiency shows, for a simple MBPP task, on a A100 GPU server to prepare for 374 instances takes about more than 18 gpu hours to complete the process of Self-Output to generate trainable data. While STM only requires 4 minutes to calculate token perplexity.
Table 11: Comparison of efficiency between Self-Output and STM in terms of gpu hours
Model Name Task Name Number of Sample Overall Correctness GPU hours
Ground Truth MBPP 374 100% 0
STM Llama3-8B Insruct MBPP 374 100% 4.5 minutes
Self-Output Llama3-8B Instruct MBPP 213 57% 16.8 hours
Rephrase Llama3-8B Instruct MBPP 327 87.4% 36.8 minutes
C.2 Model Training Resource
We train meta-llama/Meta-Llama-3-8B-Instruct, unsloth/mistral-7b-instruct-v0.3 and google/gemma-2-2b-it using two NVIDIA A100 Tensor Core GPU with VRAM 40GB for each GPU and additional RAM of 96GB. In terms of gpu hour, Our training experiments takes at most 2 hours for MBPP task, 4 to 6 hours for MATH task. As for the evaluation experiments for other out-domain tasks, evaluation in ARC-Challenge takes 3.5 hours, evaluation in GSM8k takes 4.5 hours, and evaluation in BIRD task takes 7.5 hours.
C.3 Evaluation Prompt
For each task task, we design different prompts or directly apply the prompts from the dataset. Since we use regular expression and LLM to parse out the answer we need to evaluate, so besides the task instructions, we also add different answer format instructions for different tasks for the inference prompt to parse model answer correctly.
In MBPP task, we combine a problem description (‘text‘), an answer format instruction and testing cases (‘test_list‘) into the user prompt. The generated codes will be executed with the test case to match the execution results of ground truth codes. The following TextBox demonstrate an example of MBPP data prompts:
MBPP prompt user prompt Please refer the given test cases and generate a python function for my problem. Make sure the written code is wrapped in code block : ```python $<yourcode>$``` >>>much-greater-thanabsent>>>> > > Problem:{text}>>>much-greater-thanabsent>>>> > > Test Cases:{test_list}
In MATH task, we add a prefix instruction before the problem description (‘problem‘) as the inference prompt. The answer will be parsed from the generated output using format of ‘\$ANSWER‘. The following TextBox demonstrate and example of MATH data prompts:
MATH prompt Solve the following math problem step by step. The last line of your response should be of the form Answer: \$ANSWER (without quotes) where \$ANSWER is the answer to the problem. problem Remember to put your answer on its own line after "Answer:", and you do not need to use a boxed command.
For OOD task,
In ARC-Challenge task, we add a Question-Answering task instruction and an answer format instruction before the ‘question‘ and answer choices (‘choices‘) as the user role prompt, and directly use the ‘answerKey‘ value to match with the parsed results from the output. The TextBox shows the user prompt is like:
ARC Challenge prompt user prompt Answer the following multiple choice question. The last line of your response should be of the following format: ’Answer: \$$\$LETTER^{\prime}$ (without quotes) where LETTER is one of ABCD. Question: {question} Choices: {choices}
In GSM8k task, we add a Question-Answering task instruction and an answer format instruction before the ’question’ as the user role prompt. To evaluate the output, we directly parse the format of ’####$\#\#\#\#<number\,\,only\,\,answer>$’ from the generated results, and math it with the ’answer’ value from the dataset. The TextBox shows the user prompt is like:
GSM8k prompt user prompt You are given a grade school math question. Please answer the question in the following format: Q: $<$question$>$ A: $<$Think step by step here$>$ #### $<$number only answer$>$ Format requirements : you must first output your reasoning before finalized with the " #### " format followed by the final numeric answer.
In BIRD task, we basically follow the same prompt as [32] does. A database schema (schema) is followed by a text-to-sql instruction and the question query (question) as the Textbox below shows:
BIRD prompt user prompt {schema} -- Using valid SQLite, answer the following question for the tables provided above.
-- Question: {question}
Now, generate the correct SQL code directly (Do NOT generate other text except the SQL code):
Appendix D Other STM alternatives
Besides cross-scale setting, inspired by [15, 17], we also investigate possible STM settings that acquire a model trained on ground truth data: Differential Perplexity Filtering (DPF): This approach leverages the perplexity changes induced by model training on unfiltered ground truth data. By computing the perplexity differential between the base and fine-tuned models, we identify tokens that demonstrate improved learnability during unrestricted training. The final training set is then constructed with tokens that exhibited reduced perplexity scores after fine-tuning, which potentially captures naturally learnable patterns in the data.
STM with DPF: This method extends the Differential Perplexity Filtering by incorporating our STM threshold-based masking mechanism. By combining both approaches, we aim to benefit from both the identification of learnable tokens through SFT and the prevention of high-perplexity token influence through STM, potentially offering a more robust token selection strategy than either method alone.
Table 12 shows the results for the best-performance models of the original STM and the other alternative of STM. We found that using DPF directly or applying STM on DPF would mask a much higher rate of tokens in training data, which leads to a worse training quality in terms of in-domain and out-of-domain tasks performances. One setting different from the previous works is that we did not prepare a high quality or high scored data for the reference model to train. Instead, we use ground truth data to filter the "unlearnable" tokens out. However, even though we mask out unlearnable tokens out, the remaining unfiltered tokens are still too few to maintain good quality of training.
Table 12: List of alternative token selection methodology supervised fine-tuned on MATH in Gemma2-2B Instruct. Here we choose Gemma2 series because Gemma2 has feasible model sizes to implement the cross-scale setting for comparison.
Gemma2-2B Filter MATH MBPP GSM8K ARC BIRD
Base model 0% 28.1 51.1 58.5 74.7 13.9
Ground 0% 21.7 45.8 19.0 29.5 5.1
Truth
$STM$ 23.8% 30.3 49.5 55.4 75.3 13.3
${\tau=2.5}$
${STM}_{9B}$ 21.4% 27.2 47.6 42.0 72.4 14.1
${\tau=2.0}$
Appendix E Correctness experiment with other models
We have also tested the ablation test of correctness with Gemma2-2B-Instruct model as Figure 13 shows. It holds the same conclusion that correctness is relevant to in-domain performance only, while the low perplexity trait of Self-Output data is more relevant to out-of-domain performance than correctness. However in Gemma2-2B setting, the number of training instance with a least one positive and one negative responses at the same time is much fewer than ground truth data and Llama’s correctness data although it has more trainable tokens due to different tokenizers are applied. That is to say, to maintain the same training data amount between each correctness rate dataset, we have to choose a smaller training data size for Gemma2-2B. Therefore, the effectiveness of data correctness is not as obvious as Llama3-8B model does.
Table 13: The performances of models trained with different correctness rate of MBPP Self-Output dataset and trainable tokens number. Values of a cell block in gray represents the in-domain testing performance of the Gemma2-2B Instruct trained with in-domain task data.
rejection rate Trained tokens MBPP MATH ARC GSM8k BIRD
GT 86,632 .400 .193 .229 .716 .051
Self Output
100% 87,632 .500 .287 .742 .563 .134
49% 90,594 .489 .286 .732 .564 .134
0% 90,416 .497 .284 .742 .563 .135
base model - .511 .281 .747 .585 .139
Appendix F Regularization effects on OOD
We apply a range for searching best combination of weight decay and dropout rate hyperparameters when training an MBPP task on Llama3-8B-Instruct model. The weight decay is set from 0.0 to 0.5 with interval 0.1, and dropout rate is set from 0.05 to 0.5 with an interval of 0.05 to 0.1. And we pick the top 3 best performing weight decay value with its dropout rate, and choose the best 2 performances among the checkpoints of 50 steps to 300 steps (about 5 to 6 epochs training) as our results. Our experiment results as Figure 10 show that, when applying dropout or weight decay on ground truth data, the model trained with in-domain task like MBPP still overfits on training data and the testing performance degrades. On the other hand, the out-domain performance of MATH as Figure 11 shows, we found that some of the results perform better than the original Llama in out-domain task, but their in-domain task performances are found worse than the base model’s performance. It implies that the widely adopted regularization techniques are hard to optimized for both in-domain task and out-domain task performance.
Figure 10: The in-domain performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama3-8B-Instruct) and comparison between the original Llama3-8B-Instruct model’s performance on MBPP testing dataset.
Figure 11: The out-domain MATH performance of all combinations of regularization parameters on MBPP ground-truth data model (based on Llama3-8B-Instruct) and comparison between the original Llama3-8B-Instruct model’s performance on MATH testing dataset.
Appendix G Rank analysis on LoRA weights
To further understand this behavior, we examine the relationship between weight updates and the information-theoretic properties of the training data. For a given sequence of tokens, we define the surprisal as $I(x,y)=-\sum_{t=1}^{T}\log P(y_{t}|x_{<t};\theta)$, where $P(y_{t}|x_{<t};\theta)$ represents the model’s predicted probability for the true token $y_{t}$ given the preceding context. In LoRA, weight updates for each transformer layer are parameterized as low-rank matrices $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and $r\ll\min(d,k)$. Through Singular Value Decomposition (SVD), $\Delta W=U\Sigma V^{\top}$, we can analyze both the magnitude of singular values in $\Sigma$ and its effective rank given a threshold $\tau$.
We hypothesize that training on low-surprisal data (Self-Output or Rephrase) requires smaller weight adjustments compared to high-surprisal data (Ground Truth). This is because high-surprisal sequences induce gradients that differ significantly from those generated by data the model is already familiar with, necessitating adjustments along more independent directions in the parameter space. Our empirical analysis, comparing LoRA weight updates across training conditions with identical hyperparameters and training steps, confirms this hypothesis: Ground Truth data consistently produces updates with higher effective rank, as evidenced by the singular value spectrum of $\Delta W$. This aligns with our theoretical framework, where high-surprisal data requires the model to adjust its parameters along more dimensions to minimize the loss for these unexpected sequences.
Figure 12: Effective singular value rank on layers 17th to 32th on self-attention output projection LLaMA 3 8B LoRA weights finetuned on MBPP at step 164
Figure 13: Singular values on layers 17th to 32th on self-attention output projection LoRA weights finetuned on MBPP at step 164